Freda is starting a youth group. She is the only member now, but her plans are to have
every member find 2 new members each week. If her plan works, how many
members will the group have had the end of 1 month.

Answer by ikleyn(52781)

.

To solve the problem, I make this table of three lines.


The upper line counts weeks.

Middle line shows the number of new members to join (the increment).

The lowest line shows current total members.



week                 0               1         2         3         4


new members          1               2      3*2 = 6    9*2=18    27*2=54
to join
                 Freda herself


total current
number of members    1             1+2=3     3+6=9     9+18=27   27+54=81     <<<---===  ANSWER

Solved.

-------------

From the solution,  you see that if this scheme works,  it triples the number of members each week.